OpenStudy (anonymous):

An object is initially at rest and starts to accelerate uniformly at 3 m/s2 at time t = 0 s. At time t = 8 s, what would be the distance travelled by the object?

OpenStudy (anonymous):

s = 0.5 x 3 x 8^2 = 96 m

OpenStudy (anonymous):

May i know how you get the answer?

OpenStudy (anonymous):

X= x0 + v0t+ 1/2 a t^2

OpenStudy (anonymous):

s = 0.5aT^2

OpenStudy (anonymous):

may i know why do you need the 0.5?

OpenStudy (anonymous):

1/2 a t^2

OpenStudy (anonymous):

its part of the formula..shall i derive it??

OpenStudy (anonymous):

Please. What forumla did you use ?

OpenStudy (anonymous):

see do u know v = u + at

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No . I dont.

OpenStudy (anonymous):

ohk dx/dt = v isnt it?

OpenStudy (anonymous):

no :P

OpenStudy (anonymous):

Oh yeah. I think i know how you got the first one. You re -arranged the acceleration formula ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so v = u + at dx/dt = u + at dx = (u+at)dt

OpenStudy (anonymous):

now integrate both sides

OpenStudy (anonymous):

wt do u get?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

I never learn integrate ?

OpenStudy (anonymous):

you dont need the displacement function anyway newbs

OpenStudy (anonymous):

ohk never mind u get x = ut + 0.5 at^2 ull understand when u learn integration

OpenStudy (anonymous):

v= 3t since it is always moving in the same direction distance =\[\int\limits_{0}^{8} 3t dt \]

OpenStudy (anonymous):

= (3/2) 8^2

OpenStudy (anonymous):

3 x 32 = 96m